Officer Bob issued 41 parking tickets that totaled $655.00. How many $15 tickets and how many $20 tickets did he issue?
Answer by tutorcecilia(2152)

Let x=$15
Let y=$20
x+y= 41
so, y=41-x
.
15x+20y= $655.00
15x+20(41-x)=655
15x+820-20x=655
-5x=655-820
-5x=-165
x=33
.
y=41-x
so, y=41-33=8